OpenStudy (christos):

Solve 4^x -2^x - 12 = 0

sam (.sam.):

4^x -2^x - 12 = 0 2^(2x)-2^x-12=0 let 2^x=u u^2-u-12=0 Solve

OpenStudy (anonymous):

an alternative method could be just dividing everything by 2 and getting \[2x^2-x-6=0\]

OpenStudy (anonymous):

oh ha sorry I read the q wrong :)

OpenStudy (anonymous):

x = 2, is the only solution.